OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (calculusxy):

I see that you are posting these types of questions. Do you know how to solve them?

OpenStudy (anonymous):

no

OpenStudy (calculusxy):

One of my good ways would be to actually plug in the numbers

OpenStudy (anonymous):

how do i do that ?

OpenStudy (calculusxy):

I will teach that to you. Do you know what the absolute value symbols mean?

OpenStudy (anonymous):

that whats inside it is positive ?

OpenStudy (calculusxy):

Yes or, in other words, the distance from that number and 0 (that's why it's positive).

OpenStudy (calculusxy):

So what will happen to the negative numbers once you plug them inside of the absolute value?

OpenStudy (anonymous):

they will become positive

OpenStudy (calculusxy):

Good.

OpenStudy (calculusxy):

Now let's work with the first option.

OpenStudy (anonymous):

okay

OpenStudy (calculusxy):

Actually for the sake of time, let's do the second one.

OpenStudy (anonymous):

haha alright

OpenStudy (calculusxy):

If we first plug in -3 for x. \[-3|-3+7|\] First, we need to work with the numbers inside of the parenthesis. what is -3 + 7?

OpenStudy (calculusxy):

sorry not the parenthesis, i meant the absolute value symbols.

OpenStudy (calculusxy):

in other words, what is the absolute value of -3 + 7?

OpenStudy (calculusxy):

did you get the answer?

OpenStudy (anonymous):

how would i get the absolute value of that?

OpenStudy (welshfella):

x: −3|x + 7| = −12 divide both sides by -3 |x + 7| = 4 now the left side is positive but the x + 7 can be negative or positive so we have x + 7 = 4 or x + 7 = -4 solve these 2 equations

OpenStudy (calculusxy):

-3 + 7 = 4 absolute value would just be 4

OpenStudy (anonymous):

i got x = -3, x= -11

OpenStudy (welshfella):

there are 2 answers they are the solutions of x + 7 = 4 and x + 7 = -4

OpenStudy (welshfella):

i'm sure you can solve these 2 simple equations ?

OpenStudy (anonymous):

i got that answer that i posted above

OpenStudy (anonymous):

i got -3 for the first equation and -11 for the second

OpenStudy (welshfella):

right